OpenStudy (anonymous):

if i basketball player has a leap of 35 inches what is his hangtime? hang time function v=48t^2

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

you are solving \[35=48^t\] for t

OpenStudy (anonymous):

\[35=48t^2\] for t

OpenStudy (anonymous):

divide by 48, take the square root \[t^2=\frac{35}{48}\] \[t=\sqrt{\frac{35}{48}}\]

OpenStudy (anonymous):

i get about .88